An angle is 12 degrees less than one half of its supplement. Find the supplement.
Answer by oberobic(2304)

a + 12 = 1/2s
2a + 24 = s
.
a + s = 180
s = 180 -a
.
2a + 24 = 180 -a
3a = 156
a = 52
.
s = 180-52
s = 128
.
Check
.
Does a + 12 = 1/2s?
a+12 = 52 + 12 = 64
1/2s = 1/2*128 = 64
OK
.
Answer: The supplement is 128 degrees.
.
Done.